# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ F.5 Proof of Sample-Size Scaling Rule
The reward-induced complexity term in the bound is $C_{2}|\lambda|L_{\phi}/\sqrt{n}$. Since $|\lambda(n)|=\Theta(n^{-\alpha})$, the term becomes $\Theta(n^{-(\alpha+1/2)})$. To ensure it stays within $O(n^{-1/2})$, it is necessary and sufficient that $\alpha\geq 1/2$. If $\alpha>1/2$, the decay becomes $o(n^{-1/2})$, implying a more conservative schedule. ∎
@@
We recommend starting with the default value $\alpha=\tfrac{1}{2}$, and increasing $\alpha$ in settings with noisy data or small margin $\gamma$, where $G/\gamma$ becomes large. This yields a more conservative decay for improved stability.
F.6 Practical Notes on Stabilization
Logit and Gradient Clipping
Clipping the logits as $|f_{\theta}(x)_{k}|\leq M$ and the gradients as $\|\nabla_{\theta}\ell\|\leq G_{\max}$ provides effective upper bounds on the Lipschitz constant, thereby reducing the complexity terms in generalization bounds.
Calibration of Margin Threshold $\gamma$
Setting $\gamma$ to the $q\%$-th percentile (e.g., $q=5\sim 10$) of the margin distribution $m_{\theta}(x,y)$ on the training set prevents excessive penalties from the ramp/hinge surrogate, improving stability and calibration.
Combination with $|\lambda|$ Schedule
Using a power schedule with $\alpha=\tfrac{1}{2}$ is a reasonable default. In early training stages when $|\lambda|$ is large, increasing the saturation threshold of $\phi$ (e.g., $\gamma_{\phi}$ in exponential-type designs) helps prevent premature reward saturation.
Compatibility with Trust Region
Maintaining compactness of level sets—via either an explicit constraint $\|\theta\|\leq B$ or weight normalization—ensures the assumptions of Theorem 29 are stably satisfied throughout training.
Validation-based Early Stopping
Monitoring both the validation objective $\frac{1}{n}\sum H-\lambda\frac{1}{n}\sum\phi$ and 0–1 accuracy helps detect unstable behaviors in early training. If over-rewarding causes excessive margin stretching for incorrect predictions, the schedule constant should be reduced immediately.
Appendix G Stability Perspective and the Need for Diminishing $|\lambda|$
G.1 Uniform Stability Bound
We analyze the uniform stability of the ERM solution or fixed point of a sufficiently converged first-order method under the following objective:
$\hat{F}_{\lambda}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta;z_{i})\;-\;\lambda\,R(\theta),$
where $\hat{\theta}_{\lambda}$ denotes the output model. The uniform stability is defined as:
@@ G.1 Uniform Stability Bound
with $S^{(i)}$ denoting the dataset $S$ with the $i$th sample replaced by an independent copy.
@@
(i) Each loss $\ell(\cdot;z)$ is $L$-smooth and $L_{\ell}$-Lipschitz in $\theta$ under trust-region or logit clipping constraints. (ii) The regularizer $R$ is $\alpha_{R}$-strongly convex. (iii) The empirical loss $\tfrac{1}{n}\sum_{i}\ell(\cdot;z_{i})$ is $\mu$-strongly convex. Then, the full objective is $\mu_{\mathrm{eff}}=\mu-\lambda\alpha_{R}$ strongly convex, and the safe region is $\lambda<\mu/\alpha_{R}$ ensuring $\mu_{\mathrm{eff}}>0$.
Lemma 31 (Parameter shift from single data replacement).
Under assumptions (i)–(iii), we have
@@
where $G:=\sup_{\theta,z}\|\nabla_{\theta}\ell(\theta;z)\|$ is a uniform gradient bound within the trust region.
@@
Subtracting the first-order optimality conditions for $S$ and $S^{(i)}$, and applying the co-coercivity of $\mu_{\mathrm{eff}}$-strong convexity yields $\mu_{\mathrm{eff}}\|\Delta\theta\|\leq\|\Delta g\|$. The gradient difference from a single sample change is $\|\Delta g\|\leq 2G/n$, which gives the result. ∎
@@
Under assumptions (i)–(iii), the uniform stability satisfies
@@
In particular, the bound deteriorates as $\lambda\uparrow(\mu/\alpha_{R})$, and vanishes as $\lambda\downarrow 0$ or $n\uparrow\infty$.
Remark 16 (Constant specification).
For squared loss with $\|x\|\leq R_{x}$, we obtain $G\leq R_{x}(\|X\theta\|+\|y\|)/n$. For cross-entropy with clipped logits $|f_{\theta}|\leq M$, a bound $G\leq M$ is obtained. In practice, the empirical maximum gradient norm on the validation set provides a reasonable estimate for $G$.
G.2 Restoring $\beta\to 0$ via Scheduling
Proposition 33 (Stability under $|\lambda(n)|$ scheduling).
Suppose $\lambda(n)=\lambda(n_{0})\left(\tfrac{n_{0}}{n}\right)^{\alpha}$. If the safety condition $\lambda(n)<\mu/\alpha_{R}$ is maintained, then the uniform stability satisfies
$\beta(n)\ \leq\ \frac{2\,L_{\ell}\,G}{\,n\,\big{(}\mu-\lambda(n)\alpha_{R}\big{)}}\ =\ \frac{2\,L_{\ell}\,G}{\,n\mu}\ \cdot\ \frac{1}{1-\tfrac{\alpha_{R}}{\mu}\lambda(n)}\ \xrightarrow[n\to\infty]{}0.$
In particular, if $\alpha>0$, then $\beta(n)=\tfrac{2L_{\ell}G}{n\mu}(1+o(1))$, recovering the $O(1/n)$ rate of stability.
@@ G.2 Restoring $\beta\to 0$ via Scheduling
From a stability standpoint, any $\alpha>0$ suffices. However, to balance other aspects discussed throughout the paper — such as the $|\lambda|/\sqrt{n}$ term in generalization bounds and degrees-of-freedom control in regression — we recommend $\alpha\geq 1$ for regression and $\alpha\geq\tfrac{1}{2}$ for classification.
G.3 Checklist
Estimating the safety region
@@ G.3 Checklist
Set an initial value $\lambda_{0}$ at sample size $n_{0}$, and schedule $\lambda(n)=\lambda_{0}(n_{0}/n)^{\alpha}$. Recommended values are $\alpha\geq 1$ for regression and $\alpha\geq\tfrac{1}{2}$ for classification. The variable $n$ may also refer to cumulative sample size under batch-wise or epoch-wise training.
@@
At each epoch, log the maximum gradient norm $\max_{i}\|\nabla\ell(\theta;z_{i})\|$ and the effective strong convexity $\mu_{\mathrm{eff}}=\mu-\lambda\alpha_{R}$, then monitor estimates of $\tfrac{L_{\ell}G}{n\mu_{\mathrm{eff}}}$ accordingly.
Clipping and trust regions
Apply logit/gradient clipping or a norm constraint $\|\theta\|\leq B$ to bound $G$ and $L_{\ell}$. This helps reduce the constants in Theorem 32.
Avoiding degeneracy
If the effective curvature $\mu_{\mathrm{eff}}$ falls below a threshold $\mu_{\min}=\eta_{\mu}\mu$, where $\eta_{\mu}\in[0.1,0.5]$, then decay $\lambda$ immediately or terminate training.
Weighted regularization $W\succeq 0$
If $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$, then replace $\alpha_{R}$ with $\sigma_{\min}(W)$. The safety condition becomes $\lambda<\mu/\sigma_{\min}(W)$. If instability is observed along low-spectrum directions, increase the spectral mass of $W$ to raise $\alpha_{R}$.
G.4 Matrix calculus and auxiliary identities
This section collects frequently used identities and inequalities throughout the main text and appendix. Unless otherwise specified, all matrices are real symmetric, and $\|\cdot\|$ denotes the spectral norm.
@@ G.4 Matrix calculus and auxiliary identities
For an invertible matrix $M(\lambda)$, the following hold:
@@
For invertible $A$ and appropriate shapes of $U,C,V$,
@@
In particular, if $C=\alpha I$ and $V=U^{\top}$,
$(A+\alpha UU^{\top})^{-1}=A^{-1}-A^{-1}U\big{(}\alpha^{-1}I+U^{\top}A^{-1}U\big{)}^{-1}U^{\top}A^{-1}.$
@@
Also, if $\lambda<\sigma_{\min}(A)$, then
@@
For invertible $A$ and any $B$, $\operatorname{tr}(A^{-1}B)=\operatorname{tr}(BA^{-1})$. If $0\preceq B\preceq C$ and $M\succeq 0$, then $\operatorname{tr}(MB)\leq\operatorname{tr}(MC)$.
Lemma 39 (Smoother degrees of freedom derivative).
@@
$\operatorname{tr}\big{(}X(X^{\top}X-n\lambda I)^{-1}X^{\top}\big{)}=\sum_{j}\frac{\sigma_{j}}{\sigma_{j}-\lambda},\qquad\frac{\partial}{\partial\lambda}\operatorname{tr}(\cdot)=\sum_{j}\frac{\sigma_{j}}{(\sigma_{j}-\lambda)^{2}}>0.$
@@
For a function $F$ that is $\mu$-strongly convex and $L$-smooth, the following hold:
@@
The combination of $\mu$ and $L$ repeatedly appears in stability analysis.
G.5 Variants of Noise Assumptions
@@
For general noise covariance $\Sigma_{\varepsilon}=\operatorname{Cov}(\varepsilon)\succeq 0$, the optimism identity for a linear smoother $\hat{f}=Sy$ generalizes to
@@
When $\Sigma_{\varepsilon}=\tau^{2}I$, this reduces to the classical form $\tfrac{2\tau^{2}}{n}\operatorname{tr}(S)$. Under diagonal heteroscedasticity $\Sigma_{\varepsilon}=\operatorname{diag}(\tau_{i}^{2})$, the term becomes $\sum_{i}\tau_{i}^{2}S_{ii}$, where higher leverage points contribute more to the variance term.
@@
If each $\varepsilon_{i}$ is sub-Gaussian with parameter $\sigma$, the train–test gap still concentrates via Bernstein-type inequalities. With probability at least $1-\delta$,
@@
where the constant $C$ may depend on $\|S\|_{F}$ or $\|S\|\,\operatorname{tr}(S)$. Hence, the $\sqrt{\log(1/\delta)/n}$ term in the generalization bound of the main text remains valid in order.
@@
When the sub-Gaussian assumption fails, using robust losses such as the Huber loss or median-of-means (MOM) regression approximately preserves the optimism identity. The deviation bound can still be restored to $\tilde{O}(\sqrt{1/n})$. In practice, one may estimate the contamination ratio $q$ and set the Huber threshold $\kappa$ to the corresponding quantile of the marginal distribution.
@@
Under heteroscedasticity, weighted least squares with $W_{\varepsilon}\approx\Sigma_{\varepsilon}^{-1}$ yields the modified smoother
@@
and the variance term in the optimism identity becomes
@@
under the approximation $W_{\varepsilon}\Sigma_{\varepsilon}\approx I$.
G.6 Counterexamples and Failure Cases
Deviating from recommended power schedules
In regression, if the regularization strength $|\lambda(n)|\propto n^{-\alpha}$ is not followed and $\alpha<0$ (i.e., $|\lambda|\uparrow$), the effective degrees of freedom $\operatorname{tr}(S_{\lambda})$ increase with $n$. As a result, the variance penalty in the optimism identity
@@ G.6 Counterexamples and Failure Cases
Failure in spectral weighting design
Consider $\hat{\Sigma}=\operatorname{diag}(1,\delta)$ for $0<\delta\ll 1$, and set $W=\operatorname{diag}(1,\delta^{-q})$ for $q>0$. Then,
@@
The denominator in the low-variance direction becomes sharply small ($\delta^{1+q}-\lambda$), forcing an overly small admissible $\lambda$. Even slight changes in $\lambda$ can cause wild fluctuations in the degrees of freedom. This illustrates a classic failure case where excessively large $w_{j}$ are assigned to small $\sigma_{j}$ directions.
Unbounded loss in classification
For separable data, a negative $L_{2}$ penalty $\tfrac{\lambda}{2}\|\theta\|^{2}$ leads to diverging loss: $H\to 0$ while $\tfrac{\lambda}{2}t^{2}\to-\infty$ along $\theta=tu$, hence $\hat{F}_{\lambda}\to$-$\infty$ (see Appendix F.1). Using only negative $L_{2}$ regularization in classification without trust region or bounded margin rewards is strongly discouraged.
Inconsistent initialization for DOF targeting
To solve $\tfrac{1}{n}\operatorname{tr}(S_{\lambda})=\kappa$, the small-$\lambda$ approximation
$\lambda^{(0)}\approx\frac{\kappa n-r}{\sum_{j}1/\sigma_{j}}$
may yield negative values. Using such initial guesses can cause divergence in the root-finding process. In this case, it is safer to clip $\lambda^{(0)}\leftarrow 0$ and use bisection to find a unique solution.
Insufficient margin for stability
From a stability perspective, if $\mu_{\mathrm{eff}}=\mu-\lambda\alpha_{R}$ becomes too small, the bound $\beta=O(1/(n\mu_{\mathrm{eff}}))$ worsens. In practice, it is recommended to enforce $\lambda\leq(1-\epsilon)\mu/\alpha_{R}$, and to decay $\lambda$ immediately once $\mu_{\mathrm{eff}}$ drops below a predefined threshold.

### v2

@@ G.2 Proof of the sample-size scaling rule
The reward-induced complexity term in the bound scales as $C_{2}|\lambda|L_{\phi}/\sqrt{n}$. Since $|\lambda(n)|=\Theta(n^{-\alpha})$, this term is $\Theta(n^{-(\alpha+1/2)})$. To keep it at most $O(n^{-1/2})$, it is necessary and sufficient that $\alpha\geq 1/2$. If $\alpha>1/2$, the term is $o(n^{-1/2})$, yielding a more conservative decay. ∎
@@
Use $\alpha=\tfrac{1}{2}$ as a default. If the data are noisy or the margin width $\gamma$ is small so that $G/\gamma$ is large, choose a larger $\alpha$ to enforce a more conservative decay.
G.3 Practical notes on stabilization
Logit/gradient clipping.
Clipping logits by $|f_{\theta}(x)_{k}|\leq M$ and gradients by $\|\nabla_{\theta}\ell\|\leq G_{\max}$ provides effective upper bounds on Lipschitz constants and reduces the complexity terms.
Calibrating the margin threshold $\gamma$.
Set $\gamma$ to a lower $q\%$ quantile of the training margin distribution $m_{\theta}(x,y)$ to avoid excessive penalization by the ramp/hinge surrogate (e.g., $q=5\!\sim\!10$).
Combining with the $|\lambda|$ schedule.
Adopt the $\alpha=\tfrac{1}{2}$ power-law decay as a default; when $|\lambda|$ is large at early stages, increase the saturation threshold of $\phi$ (e.g., the $\gamma_{\phi}$ in the exponential design) to prevent premature reward saturation.
Harmony with trust regions.
Maintain compact level sets via $\|\theta\|\leq B$ or Weight Normalization so that the assumptions of Theorem 30 are reliably satisfied.
Validation-driven stopping.
Monitor both the empirical objective $\tfrac{1}{n}\sum H-\lambda\tfrac{1}{n}\sum\phi$ and a 0–1 metric. If early training exhibits signals of over-rewarded margin expansion (with large $|\lambda|$), immediately reduce the schedule constant.
Appendix H Stability perspective and the necessity of decreasing $|\lambda|$
H.1 Uniform stability bound derivation
In this section we analyze uniform stability when the empirical risk is augmented with a negative reward term:
$\hat{F}_{\lambda}(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta;z_{i})\;-\;\lambda\,R(\theta).$
Let $\hat{\theta}_{\lambda}$ denote the ERM solution or the fixed point of a sufficiently converged first-order method. The uniform stability is defined as
@@ H.1 Uniform stability bound derivation
where $S^{(i)}$ denotes the dataset obtained from $S$ by replacing its $i$-th sample with an independent draw from the same distribution.
@@
(i) Each $\ell(\cdot;z)$ is $L$-smooth, and under a trust-region or logit clipping, $\ell$ is $L_{\ell}$-Lipschitz in $\theta$. (ii) The reward term $R$ is $\alpha_{R}$-strongly convex. (iii) The empirical loss $\tfrac{1}{n}\sum_{i}\ell(\cdot;z_{i})$ is $\mu$-strongly convex. Then the full objective is $\mu_{\mathrm{eff}}=\mu-\lambda\alpha_{R}$ strongly convex, and under the safety condition $\lambda<\mu/\alpha_{R}$ we have $\mu_{\mathrm{eff}}>0$.
Lemma 32 (Parameter shift under one-sample replacement).
Under assumptions (i)–(iii),
@@
where $G:=\sup_{\theta,z}\|\nabla_{\theta}\ell(\theta;z)\|$ is a gradient bound valid within the trust region.
@@
Subtracting the first-order optimality conditions for $S$ and $S^{(i)}$ and applying co-coercivity from $\mu_{\mathrm{eff}}$-strong convexity gives $\mu_{\mathrm{eff}}\|\Delta\theta\|\leq\|\Delta g\|$. The gradient perturbation caused by replacing one sample satisfies $\|\Delta g\|\leq 2G/n$, yielding the result. ∎
@@
Under assumptions (i)–(iii),
@@
In particular, as $\lambda$ approaches $\mu/\alpha_{R}$, the bound deteriorates. Conversely, as $\lambda\to 0$ or $n\to\infty$, the stability parameter $\beta$ converges to zero.
Remark 20 (Specification of constants).
For the squared loss with bounded covariates $\|x\|\leq R_{x}$, one can bound $G\leq R_{x}(\|X\theta\|+\|y\|)/n$. For the cross-entropy loss with logit clipping $|f_{\theta}|\leq M$, a typical bound is $G\leq M$. In practice, $G$ can be estimated by the maximum gradient observed on the validation set.
H.2 Recovering $\beta\to 0$ with schedules
Proposition 34 ($|\lambda(n)|$ schedule and stability).
Let $\lambda(n)=\lambda(n_{0})\big{(}\tfrac{n_{0}}{n}\big{)}^{\alpha}$. If the safety condition $\lambda(n)<\mu/\alpha_{R}$ is maintained, then
$\beta(n)\ \leq\ \frac{2\,L_{\ell}\,G}{\,n\,\big{(}\mu-\lambda(n)\alpha_{R}\big{)}}\ =\ \frac{2\,L_{\ell}\,G}{\,n\mu}\ \frac{1}{1-\tfrac{\alpha_{R}}{\mu}\lambda(n)}\ \xrightarrow[n\to\infty]{}\ 0.$
In particular, if $\alpha>0$, then $\beta(n)=\tfrac{2L_{\ell}G}{n\mu}\big{(}1+o(1)\big{)}$, thus recovering the canonical $1/n$ rate of uniform stability.
@@ H.2 Recovering $\beta\to 0$ with schedules
For stability alone, any $\alpha>0$ suffices. However, when combined with the $|\lambda|/\sqrt{n}$ term arising in the generalization bound and the degree-of-freedom control in regression, we recommend $\alpha\geq 1$ for regression and $\alpha\geq\tfrac{1}{2}$ for classification.
H.3 Checklist
Safety interval estimation
@@ H.3 Checklist
Initialize $\lambda_{0}$ at $n_{0}$, and schedule $\lambda(n)=\lambda_{0}(n_{0}/n)^{\alpha}$. Recommended $\alpha$: at least $1$ for regression and at least $\tfrac{1}{2}$ for classification. It is acceptable to interpret $n$ as the cumulative number of samples with increasing batch size or epoch progression.
@@
At each epoch, log the maximum gradient norm over random samples, $\max_{i}\|\nabla\ell(\theta;z_{i})\|$, and the effective strong convexity $\mu_{\mathrm{eff}}=\mu-\lambda\alpha_{R}$. Monitor the estimate of $\tfrac{L_{\ell}G}{n\mu_{\mathrm{eff}}}$ as a stability diagnostic.
Clipping and trust region
Apply logit/gradient clipping or impose a constraint $\|\theta\|\leq B$ to secure effective upper bounds on $G$ and $L_{\ell}$. This reduces the constants in Theorem 33.
Boundary avoidance
If $\mu_{\mathrm{eff}}$ drops below a threshold $\mu_{\min}$ (e.g., $\mu_{\min}=\eta_{\mu}\,\mu$, with $\eta_{\mu}\in[0.1,0.5]$), immediately decay $\lambda$ or halt training.
Weighted reward $W\succeq 0$
For $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$, replace $\alpha_{R}$ with $\sigma_{\min}(W)$. The safety condition becomes $\lambda<\mu/\sigma_{\min}(W)$. If divergence is observed along small singular value directions, enlarge the spectrum of $W$ to increase $\alpha_{R}$.
Appendix I Auxiliary results and arithmetic facts
I.1 Matrix calculus and supporting lemmas
This section collects frequently used identities and inequalities that appear throughout the main text and appendix. Unless otherwise specified, all matrices are real symmetric, and $\|\cdot\|$ denotes the spectral norm.
@@ I.1 Matrix calculus and supporting lemmas
For invertible $M(\lambda)$,
@@
For invertible $A$ and matrices $U,C,V$ of compatible dimensions,
@@
In particular, if $C=\alpha I$, $V=U^{\top}$,
$(A+\alpha UU^{\top})^{-1}=A^{-1}-A^{-1}U(\alpha^{-1}I+U^{\top}A^{-1}U)^{-1}U^{\top}A^{-1}.$
@@
If $\lambda<\sigma_{\min}(A)$,
@@
For invertible $A$ and arbitrary $B$, $\operatorname{tr}(A^{-1}B)=\operatorname{tr}(BA^{-1})$. If $0\preceq B\preceq C$ and $M\succeq 0$, then $\operatorname{tr}(MB)\leq\operatorname{tr}(MC)$.
Lemma 40 (Derivative of smoother degrees of freedom).
@@
$\operatorname{tr}\big{(}X(X^{\top}X-n\lambda I)^{-1}X^{\top}\big{)}=\sum_{j}\frac{\sigma_{j}}{\sigma_{j}-\lambda},\quad\frac{\partial}{\partial\lambda}\operatorname{tr}(\cdot)=\sum_{j}\frac{\sigma_{j}}{(\sigma_{j}-\lambda)^{2}}>0.$
@@
For a $\mu$-strongly convex and $L$-smooth function $F$,
@@
These properties are repeatedly invoked in the stability analysis when combining $\mu$ and $L$.
I.2 Variants of noise assumptions
@@
For a general covariance $\Sigma_{\varepsilon}=\operatorname{Cov}(\varepsilon)\succeq 0$, the optimism identity of a linear smoother $\hat{f}=Sy$ becomes
@@
With homoscedastic noise $\Sigma_{\varepsilon}=\tau^{2}I$, this reduces to $\tfrac{2\tau^{2}}{n}\operatorname{tr}(S)$. For diagonal heteroscedastic noise $\Sigma_{\varepsilon}=\operatorname{diag}(\tau_{i}^{2})$, we obtain $\operatorname{tr}(S\Sigma_{\varepsilon})=\sum_{i}\tau_{i}^{2}S_{ii}$, so high-leverage points contribute larger variance.
@@
If each $\varepsilon_{i}$ is sub-Gaussian with parameter $\sigma$, the train–test gap concentrates in a Bernstein-type bound, yielding with probability $1-\delta$,
@@
where the constant $C$ may depend on $\|S\|_{F}$ or $\|S\|\,\operatorname{tr}(S)$. Thus the $\sqrt{\log(1/\delta)/n}$ term in the main generalization bound remains of the same order.
@@
If the sub-Gaussian assumption fails, Huber loss or median-of-means (MOM) regression can be used instead, and the optimism identity remains approximately valid. Concentration terms can be restored to $\tilde{O}(\sqrt{1/n})$. In practice, one may estimate the outlier fraction $q$ and set the Huber threshold $\kappa$ at the $q$-quantile level.
@@
In heteroscedastic settings, weighted least squares with $W_{\varepsilon}\approx\Sigma_{\varepsilon}^{-1}$ yields
@@
with the variance term in the optimism identity simplifying to
@@
which assumes $W_{\varepsilon}\Sigma_{\varepsilon}\approx I$.
I.3 Collection of counterexamples
Deterioration from violating the power schedule
If one chooses $\alpha<0$ in regression, i.e. $|\lambda|$ increasing rather than decreasing as recommended $|\lambda(n)|\propto n^{-\alpha}$, then $\operatorname{tr}(S_{\lambda})$ grows with $n$. Consequently,
@@ I.3 Collection of counterexamples
Failure of spectral weighting design
Consider $\hat{\Sigma}=\operatorname{diag}(1,\delta)$ with $0<\delta\ll 1$ and choose $W=\operatorname{diag}(1,\delta^{-q})$ for some $q>0$. Then
@@
The denominator of the small-eigenvalue direction becomes $\delta^{1+q}-\lambda$, forcing the safety boundary $\lambda<\delta^{1+q}$ to be extremely small. Even tiny changes in $\lambda$ cause violent oscillations of the degrees of freedom. This illustrates the typical failure of assigning large $w_{j}$ to small $\sigma_{j}$.
Reappearance of unboundedness in classification
For separable data with negative $L_{2}$ reward $\tfrac{\lambda}{2}\|\theta\|^{2}$, along the path $\theta=tu$ we have $H\to 0$ but $\tfrac{\lambda}{2}t^{2}\to-\infty$, hence $\hat{F}_{\lambda}\to$-$\infty$ (see Appendix F.1). Thus using negative $L_{2}$ alone in classification without a trust region or bounded margin reward is unsafe.
Inconsistent initialization for degree targeting
In solving $\tfrac{1}{n}\operatorname{tr}(S_{\lambda})=\kappa$, the small-$\lambda$ approximation $\lambda^{(0)}\approx\tfrac{\kappa n-r}{\sum_{j}1/\sigma_{j}}$ may be negative. Using such an initialization directly may prevent convergence of the monotone equation solver and lead to divergence. A safe remedy is to clip $\lambda^{(0)}\leftarrow 0$ and use bisection to locate the unique solution.
Insufficient stability margin
When the effective strong convexity $\mu_{\mathrm{eff}}=\mu-\lambda\alpha_{R}$ becomes too small, the stability bound $\beta=O(1/(n\mu_{\mathrm{eff}}))$ deteriorates. In practice, one enforces $\lambda\leq(1-\epsilon)\mu/\alpha_{R}$, and if $\mu_{\mathrm{eff}}$ falls below a threshold, immediately decays $\lambda$.